having troble helping my son with the following eq:
F(x)=mx+b /
The work sheet gives 3, 5 and 7 as inputs (x) and resulting output (F) being 5, 7, 9
F(x) = x+2
They want him to take it out two more places.
I know the answer.
How do I explain this to him.

Answer by Edwin McCravy(20086)

having troble helping my son with the following eq:
F(x)=mx+b /
The work sheet gives 3, 5 and 7 as inputs (x) and resulting
output (F) being 5, 7, 9
F(x) = x+2

The reason the input of 3 has the resulting output 5 is because if you
replace x by 3 in F(x) = x+2
you get:

F(3) = 3+2

F(3) = 5

The reason the input of 5 has the resulting output 7 is because if you
replace x by 5 in F(x) = x+2
you get:

F(5) = 5+2

F(5) = 7

The reason the input of 7 has the resulting output 9 is because if you
replace x by 7 in F(x) = x+2
you get:

F(7) = 7+2

F(7) = 9

Therefore to find the resulting output for the input 9, you replace
x by 9 in F(x) = x+2
and get:

F(9) = 9+2

F(9) = 11

So the input of 9 gives the resulting output of 11

Therefore to find the resulting output for the input 11, you replace
x by 11 in F(x) = x+2
and get:

F(11) = 11+2

F(11) = 13

So the input of 11 gives the resulting output of 13

I suppose this is what they mean by "take it out two more places".

Edwin